Nico is saving money for his college education he invests some money at 7% and 1300 less than that amount at 6% the investment produced a total of 182 interest in one year how much did he invested at each rate
Answer by Boreal(15235)

x at 7%
x-1300 at 6%
.07x+.06(x-1300)=182
.07x+.06x-78=182
.13x=260
x=$2000 @ 7%=$140
x-1300=$700@6%=$42
Total=$182